The life of a 60- watt light bulb in hours is known to be normally distributed with σ = 25
hours. Create 5 different random samples of 100 bulbs each which has a mean life of x_bar ~
1000 hours and perform one-way ANOVA with state it.
1
Expert's answer
2020-07-02T19:14:49-0400

The total sample size is "N=500." Therefore, the total degrees of freedom are:


"df_{total}=500-1=499"

The between-groups degrees of freedom are "df_{between}=5-1=4," and the within-groups degrees of freedom are:


"df_{within}=df_{total}-df_{between}=499-4=495"

"\displaystyle\sum_{i,j}X_{ij}=499712"

"\displaystyle\sum_{i,j}X_{ij}^2=499691630"

"SS_{total}=\displaystyle\sum_{i,j}X_{ij}^2-\dfrac{1}{N}(\displaystyle\sum_{i,j}X_{ij})^2=267464.112"

"SS_{within}=266084.42"

"SS_{between}=1379.692\n\u200b"

"MS_{between}=\dfrac{SS_{between}}{df_{between}}=\dfrac{1379.692}{4}=344.923"

"MS_{within}=\dfrac{SS_{within}}{df_{within}}=\dfrac{266084.42}{495}=537.544"

"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{344.923}{537.544}=0.642"

The following null and alternative hypotheses need to be tested:

"H_0: \mu_1=\mu_2=\mu_3=\mu_4=\mu_5"

"H_1:" Not all means are equal.

The above hypotheses will be tested using an F-ratio for a One-Way ANOVA.


Based on the information provided, the significance level is "\alpha=0.05," and the degrees of freedom are "df_1=4" and "df_2=4," therefore, the rejection region for this F-test is "R=\{F:F>F_c=2.39\}."

Test Statistics


"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{344.923}{537.544}=0.642"

Since it is observed that "F=0.642<2.39=F_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 5 population means are equal, at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value is "p=0.633," and since "p=0.633\geq0.05,"

it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 5 population means are equal, at the "\alpha=0.05" significance level.